OpenStudy (gabylovesyou):

What is the slope of a line that passes through the point (−5, 3) and is parallel to a line that passes through (2, 13) and (−4, −11)?

ganeshie8 (ganeshie8):

parallel lines will have equal slopes.

ganeshie8 (ganeshie8):

so just compute the slope of line passing thru : (2, 13) and (−4, −11)

OpenStudy (whpalmer4):

Here the point that your line goes through is unnecessary information, added to possibly confuse you :-)

OpenStudy (gabylovesyou):

i put it in geogebra, thats what u mean ? @ganeshie8 and ok @whpalmer4 thanks

OpenStudy (whpalmer4):

slope of the line can be found with \[m = \frac{(y_2-y_1)}{(x_2-x_1)}\]where \((x_1,y_1)\) and \((x_2,y_2)\) are your points

ganeshie8 (ganeshie8):

we need to use slope formula for this.. (2, 13) and (−4, −11) slope = diff in y / diff in x = -11-13 / -4-2 = -24/-6 = ?

OpenStudy (gabylovesyou):

m = 4

ganeshie8 (ganeshie8):

yes that easy it is. im sure the other extra point confused u... :)

OpenStudy (gabylovesyou):

yes it did.. lol

OpenStudy (whpalmer4):

many problems would ask you to put a parallel line through that other point, do you know how to do so?

OpenStudy (gabylovesyou):

no clue lol

OpenStudy (whpalmer4):

Okay, here's the formula you might use: \[y-y_0 = m(x-x_0)\] which is the equation for a line with slope \(m\) passing through point \((x_0,y_0)\)

OpenStudy (whpalmer4):

So for the point they gave us, and the slope you found, the equation would be \[y-3 = 4(x-(-5))\]\[y-3=4(x+5)\]

OpenStudy (gabylovesyou):

ohh ok